# SimpleMem: Efficient Lifelong Memory for LLM Agents

### v2

@@ Abstract
To support long-term interaction in complex environments, LLM agents require memory systems that manage historical experiences. Existing approaches either retain full interaction histories via passive context extension, leading to substantial redundancy, or rely on iterative reasoning to filter noise, incurring high token costs. To address this challenge, we introduce SimpleMem, an efficient memory framework based on semantic lossless compression. We propose a three-stage pipeline designed to maximize information density and token utilization: (1) Semantic Structured Compression, which distills unstructured interactions into compact, multi-view indexed memory units; (2) Online Semantic Synthesis, an intra-session process that instantly integrates related context into unified abstract representations to eliminate redundancy; and (3) Intent-Aware Retrieval Planning, which infers search intent to dynamically determine retrieval scope and construct precise context efficiently. Experiments on benchmark datasets show that our method consistently outperforms baseline approaches in accuracy, retrieval efficiency, and inference cost, achieving an average F1 improvement of 26.4% while reducing inference-time token consumption by up to 30×, demonstrating a superior balance between performance and efficiency. Code is available at https://github.com/aiming-lab/SimpleMem.
@@ 2.1 Semantic Structured Compression
Specifically, first, incoming dialogue is segmented into overlapping sliding windows $W$ of fixed length, where each window represents a short contiguous span of recent interaction. These windows serve as the basic units for processing. Unlike traditional approaches that rely on rigid heuristic filters, we employ a Semantic Density Gating mechanism. We model the information assessment as an implicit reasoning task performed by the foundation model itself. The system leverages the attention mechanism of the LLM $f$ to estimate the information gain of the current window relative to the immediate history $H$.
Formally, we define the gating function $\Phi_{\text{gate}}$ as an indicator function that determines whether the semantic utility of $W$ exceeds the model’s implicit retention threshold $\tau$:
$\Phi_{\text{gate}}(W)=\mathbb{I}\left(f(W\mid H)>\tau\right).$ (1)
Windows where $\Phi_{\text{gate}}(W)=0$ are low-density (e.g., pure chitchat) and are discarded efficiently. This soft gating approach allows the system to capture subtle semantic nuances that rigid keyword or cosine-based filters often miss.
For windows that pass the gate ($\Phi_{\text{gate}}(W)=1$), the system proceeds a normalization and decomposition step. We apply a transformation $\mathcal{F}_{\text{norm}}$ to decompose the window into a set of context-independent memory units $\{m_{k}\}_{k=1}^{K}$. This transformation resolves dependencies implicit in conversational flow by converting entangled dialogue into self-contained factual statements. Here, $\mathcal{F}_{\text{norm}}$ encapsulates three operations: fact extraction ($g_{\text{ext}}$), coreference resolution ($g_{\text{coref}}$), and temporal anchoring ($g_{\text{time}}$):
$m_{k}\in\mathcal{F}_{\text{norm}}(W)=g_{\text{time}}\circ g_{\text{coref}}\circ g_{ext}(W;f).$ (2)
Here, $g_{\text{coref}}$ replaces ambiguous pronouns with specific entity names (e.g., changing "He agreed" to "Bob agreed"), and $g_{\text{time}}$ converts relative temporal expressions (e.g., transforming "next Friday" to "2025-10-24") into absolute ISO-8601 timestamps. This normalization ensures that each memory unit remains interpretable and valid independent of its original conversational context. By aggregating all resulting units $m_{k}$ across sliding windows, we obtain the complete memory set $\mathcal{M}$.
@@
To support flexible and high-fidelity retrieval, each memory unit is indexed through three complementary representations. First, at the Semantic Layer, we map the entry to a dense vector space $s_{k}$ using embedding models, capturing abstract meaning to enable fuzzy matching (e.g., retrieving "latte" when querying "hot drink"). Second, the Lexical Layer utilizes an inverted index-based sparse representation. This acts as a high-dimensional sparse vector $l_{k}$ focusing on exact keyword matches and rare proper nouns, ensuring that specific entities are not diluted in dense vector space. Third, the Symbolic Layer extracts structured metadata, such as timestamps and entity types, to enable deterministic filtering logic. Formally, given a memory unit these projections form the comprehensive indexing $\mathcal{I}$:
@@ 3 Experiments
(n=133) (n=133) (n=78) (n=70) (n=56) (n=30)
GPT-4.1-mini Full-context 27.06% 30.08% 41.03% 47.14% 32.14% 60.00% 35.40%
Mem0 40.60% 50.37% 69.23% 87.14% 48.21% 63.33% 56.40%
LightMem 85.71% 47.37% 92.30% 88.57% 21.43% 76.67% 69.20%
SimpleMem 70.68% 60.92% 79.48% 85.71% 75.00% 76.67% 72.40%
GPT-4.1 Full-context 51.88% 39.10% 70.51% 65.71% 96.43% 16.67% 56.20%
Mem0 43.61% 54.89% 75.64% 54.29% 39.29% 83.33% 55.00%
LightMem 84.96% 57.89% 89.74% 87.14% 71.43% 70.00% 76.40%
SimpleMem 66.91% 81.20% 80.76% 98.57% 64.29% 76.78% 77.60%
@@ 3.1 Experimental Setup
Implementation Details. For semantic structured compression, we use a sliding window of size $W=10$. Memory indexing is implemented using LanceDB with a multi-view design: Qwen3-embedding-0.6b (1024 dimensions) for dense semantic embeddings, BM25 for sparse lexical indexing, and SQL-based metadata storage for symbolic attributes. During retrieval, we employ adaptive query-aware retrieval, where the retrieval depth is dynamically adjusted based on estimated query complexity, ranging from $k_{\min}=3$ for simple lookups to $k_{\max}=20$ for complex reasoning queries.
@@ 3.2 Results and Analysis
Results on LongMemEval-S (Table 2) further demonstrate the robustness of SimpleMem under extreme context lengths. Using the gpt-4.1-mini backbone, SimpleMem achieves the highest average accuracy of 72.40%, outperforming LightMem (69.20%) and substantially exceeding Mem0 (56.40%) and the full-context baseline (35.40%). Gains are most pronounced in the challenging Multi-Session category, where SimpleMem attains 60.92% accuracy, compared to 47.37% for LightMem and 30.08% for full-context, highlighting its effectiveness in cross-session information integration under severe context constraints.
When scaled to the more capable gpt-4.1 backbone, SimpleMem maintains its state-of-the-art performance with an average accuracy of 77.60%. It is particularly strong in the Single-Session User task (98.57%), demonstrating near-perfect recall of immediate user-provided details. While LightMem exhibits strong performance on temporally specific queries, SimpleMem offers a more balanced profile: it avoids catastrophic failures in assistant-focused recall (where LightMem drops to 21.43% on the mini model) while maintaining high accuracy in complex multi-session retrieval. This balance suggests that SimpleMem’s structured indexing strategy effectively disentangles episodic noise from salient facts, providing a reliable memory substrate for long-term interaction.
@@ 3.3 Efficiency Analysis
Table 4: Comparison of construction time, retrieval time, total experiment time, and average F1 score across different memory systems (tested on LoCoMo-10 with GPT-4.1-mini).
Model Construction Time Retrieve Time Total Time Average F1
@@ 5 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Use of AI Assistants
We acknowledge the use of AI assistants in the writing of this paper, mostly for formatting and polishing the text. All the generated content is reviewed by the authors and necessary corrections are made.

### v3

@@ Abstract
To support long-term interaction in complex environments, LLM agents require memory systems that manage historical experiences. Existing approaches either retain full interaction histories via passive context extension, leading to substantial redundancy, or rely on iterative reasoning to filter noise, incurring high token costs. To address this challenge, we introduce SimpleMem, an efficient memory framework based on semantic lossless compression. We propose a three-stage pipeline designed to maximize information density and token utilization: (1) Semantic Structured Compression, which distills unstructured interactions into compact, multi-view indexed memory units; (2) Online Semantic Synthesis, an intra-session process that instantly integrates related context into unified abstract representations to eliminate redundancy; and (3) Intent-Aware Retrieval Planning, which infers search intent to dynamically determine retrieval scope and construct precise context efficiently. Experiments on benchmark datasets show that our method consistently outperforms baseline approaches in accuracy, retrieval efficiency, and inference cost, achieving an average F1 improvement of 26.4% in LoCoMo while reducing inference-time token consumption by up to 30×, demonstrating a superior balance between performance and efficiency. Code is available at https://github.com/aiming-lab/SimpleMem.
@@ 2.1 Semantic Structured Compression
Specifically, first, incoming dialogue is segmented into overlapping sliding windows $W$ of fixed length, where each window represents a short contiguous span of recent interaction. These windows serve as the basic units for processing.
Unlike traditional approaches that rely on rigid heuristic filters or separate classification models, we employ an implicit semantic density gating mechanism integrated directly into the generation process. We model the information assessment as an instruction-following task performed by the foundation model itself. The system leverages the attention mechanism of the LLM $f$ to identify high-entropy spans within the window $W$ relative to the immediate history $H$.
Formally, we define the gating function $\Phi_{\text{gate}}$ not as a binary classifier, but as a generative filter resulting from the model’s extraction capability:
$\Phi_{\text{gate}}(W)\to\{m_{k}\}\quad\text{s.t.}\quad|\{m_{k}\}|\geq 0$ (1)
Here, the generation of an empty set ($\emptyset$) inherently signifies a low-density window (e.g., pure phatic chitchat), effectively discarding it without explicit threshold tuning. This instruction-driven gating allows the system to capture subtle semantic nuances while naturally filtering redundancy through the model’s semantic compression objectives.
For windows containing valid semantic content, the system performs a unified De-linearization Transformation $\mathcal{F}_{\theta}$. Instead of sequential independent modules, we optimize the extraction, coreference resolution, and temporal anchoring as a joint generation task. The transformation projects the raw dialogue window $W$ directly into a set of context-independent memory units $\{m_{k}\}$:
$\{m_{k}\}=\mathcal{F}_{\theta}(W;H)\approx(g_{\text{time}}\circ g_{\text{coref}}\circ g_{\text{ext}})(W).$ (2)
In this unified pass, the model follows strict instructional constraints to: (1) resolve ambiguous pronouns to specific entity names ($g_{\text{coref}}$), (2) convert relative temporal expressions into absolute ISO-8601 timestamps ($g_{\text{time}}$), and (3) atomize complex dialogue flows into self-contained factual statements. By aggregating all resulting units $m_{k}$ across sliding windows, we obtain the complete memory set $\mathcal{M}$.
@@
To support flexible and high-fidelity retrieval, each memory unit is indexed through three complementary representations. First, at the Semantic Layer, we map the entry to a dense vector space $s_{k}$ using embedding models, capturing abstract meaning to enable fuzzy matching (e.g., retrieving "latte" when querying "hot drink"). Second, the Lexical Layer utilizes an inverted index-based sparse representation. This acts as a high-dimensional sparse vector $l_{k}$ focusing on exact keyword matches and rare proper nouns, ensuring that specific entities are not diluted in dense vector space. Third, the Symbolic Layer extracts structured metadata, such as timestamps and entity types, to enable deterministic filtering logic. Formally, for a given memory unit $m_{k}$, these projections form the comprehensive indexing $\mathcal{I}$:
@@ 3 Experiments
GPT-4.1-mini Full-context 27.06% 30.08% 41.03% 47.14% 32.14% 60.00% 39.57%
Mem0 40.60% 50.37% 69.23% 87.14% 48.21% 63.33% 59.81%
LightMem 85.71% 47.37% 92.30% 88.57% 21.43% 76.67% 68.67%
SimpleMem 83.46% 60.92% 79.48% 85.71% 75.00% 76.67% 76.87%
GPT-4.1 Full-context 51.88% 39.10% 70.51% 65.71% 96.43% 16.67% 56.72%
Mem0 43.61% 54.89% 75.64% 54.29% 39.29% 83.33% 58.51%
LightMem 84.96% 57.89% 89.74% 87.14% 71.43% 70.00% 76.86%
SimpleMem 86.47% 81.20% 80.76% 98.57% 76.79% 80.00% 83.97%
@@ 3.1 Experimental Setup
Implementation Details. For semantic structured compression, we use a sliding window of size $W=20$. Memory indexing is implemented using LanceDB with a multi-view design: Qwen3-embedding-0.6b (1024 dimensions) for dense semantic embeddings, BM25 for sparse lexical indexing, and SQL-based metadata storage for symbolic attributes. During retrieval, we employ adaptive query-aware retrieval, where the retrieval depth is dynamically adjusted based on estimated query complexity, ranging from $k_{\min}=3$ for simple lookups to $k_{\max}=20$ for complex reasoning queries.
@@ 3.2 Results and Analysis
Results on LongMemEval-S (Table 2) further demonstrate the robustness of SimpleMem under extreme context lengths. Using the gpt-4.1-mini backbone, SimpleMem achieves the highest average accuracy of 76.87%, outperforming LightMem (68.67%) and substantially exceeding Mem0 (59.81%) and the full-context baseline (39.57%). Gains are most pronounced in the challenging Multi-Session category, where SimpleMem attains 60.92% accuracy, compared to 47.37% for LightMem and 30.08% for full-context, highlighting its effectiveness in cross-session information integration under severe context constraints.
When scaled to the more capable gpt-4.1 backbone, SimpleMem maintains its state-of-the-art performance with an average accuracy of 83.97%. It is particularly strong in the Single-Session User task (98.57%), demonstrating near-perfect recall of immediate user-provided details. While LightMem exhibits strong performance on temporally specific queries, SimpleMem offers a more balanced profile: it avoids catastrophic failures in assistant-focused recall (where LightMem drop significantly on the mini model) while maintaining high accuracy in complex multi-session retrieval. This balance suggests that SimpleMem’s structured indexing strategy effectively disentangles episodic noise from salient facts, providing a reliable memory substrate for long-term interaction.
@@ 3.3 Efficiency Analysis
Table 4: Comparison of construction time, retrieval time, total experiment time, and average F1 score across different memory systems (tested on LoCoMo-10 with GPT-4.1-mini; time values are reported as per-sample averages on LoCoMo-10).
Model Construction Time Retrieval Time Total Time Average F1